OpenStudy (anonymous):

Will fan and medal (: Basic geometry Two similar rectangular prisms have scale factor of 2:3, if the volume of the larger prism is 54 cm3, whats the volume of the smaller prism?

OpenStudy (anonymous):

2/3 = x/54 Cross multiply: 3x = 108 Simplify (divide by 3): x = 36 Answer: 36 cm^3

OpenStudy (anonymous):

Thank you so muchhh!

OpenStudy (anonymous):

No problem